# A unified analytical prediction for steady-state behavior of confined drop with interface viscosity under shear flow

### v3

@@ Abstract
respectively. Since they considered an unconfined pure drop ($\chi=0$ and $\text{Bq}=0$), the functions $f_{i}$ have been chosen to recover the first-order expansion in Ca of the shape-evolution equation computed by Rallison [16]: $f_{1}(\lambda)=f_{1}^{(\mbox{\tiny MM})}(\lambda)=40(\lambda+1)/[(2\lambda+3)(19\lambda+16)]$ and $f_{2}(\lambda)=f_{2}^{(\mbox{\tiny MM})}(\lambda)=5/(2\lambda+3)$. Since Rallison’s result has been recently extended by Narsimhan (N) [10] to include the effect of the interface viscosity, we matched the functions $f_{i}$ with the first-order expansion of the shape-evolution equation computed in Ref. [10], thus obtaining $f_{i}(\lambda,\text{Bq})=f_{i}^{(\mbox{\tiny N})}(\lambda,\text{Bq})$ (cfr. Eqs. (17) in Appendix .2). Notice that, in the limit of pure drop, $f_{i}^{(\mbox{\tiny N})}(\lambda,\text{Bq}=0)$ coincide with $f_{i}^{(\mbox{\tiny MM})}(\lambda)$. To include the effect of the confinement degree, we followed Minale’s approach [51], who extended $f_{i}^{(\mbox{\tiny MM})}$ to account for the effect of $\chi$. Thus, we extended $f_{i}$ as $f_{1}^{\mathrm{ext}}=f_{1}^{(\mbox{\tiny N})}(\lambda,\text{Bq})/\psi_{1}(\lambda,\hat{h},\chi)$ and $f_{2}^{\mathrm{ext}}=f_{2}^{(\mbox{\tiny N})}(\lambda,\text{Bq})\psi_{2}(\lambda,\hat{h},\chi)$, with $\psi_{i}=1+C_{s}(\hat{h})(\chi/2)^{3}\tilde{f}_{i}(\lambda)$, where the expressions for $\tilde{f}_{i}$ have to be matched with the analytical results by Shapira & Haber [24] (in Ref. [51], $f_{i}^{(\mbox{\tiny N})}$ are replaced by $f_{i}^{(\mbox{\tiny MM})}$ in the definition of $f_{i}^{\mathrm{ext}}$). To obtain an expression for $\tilde{f}_{i}$, we substituted $f_{i}$ with $f_{i}^{\mathrm{ext}}$ in Eq. (7a) and, by neglecting $\mathcal{O}(\chi^{6})$, we obtained:
@@
By comparing Eq. (8) with the first-order expansion in Ca of Eq. (6), i.e.,
$\mbox{D}(\lambda,\hat{h},\chi,\text{Ca},\text{Bq})=\text{Ca}\frac{19\lambda+16+32\text{Bq}}{16\lambda+16+32\text{Bq}}\left[1+C_{s}(\hat{h})\frac{\chi^{3}}{8}\frac{1+2.5\lambda}{1+\lambda}\right]\ ,$ (9)
it is straightforward to see that $\tilde{f}_{1}+\tilde{f}_{2}=(1+2.5\lambda)/(1+\lambda)$. This implies that, once $\tilde{f}_{2}$ is determined, $\tilde{f}_{1}$ is fixed. It is noteworthy that $\Psi(\lambda,\hat{h},\chi)$ (i.e., the term in squared parenthesis in Eq. (9)) does not depend on Bq by what we observed for the steady-state deformation (cfr. Eqs. (5) and (6) and Fig. 2). Hence, neither $\tilde{f}_{1}$ nor $\tilde{f}_{2}$ depend on Bq, and then the values computed by Minale [51] for a pure confined drop are still valid:
@@
All results hitherto discussed refer to the case of $\lambda=1$. To prove the robustness of the unified analytical expressions for both D (cfr. Eq. (6)) and $\theta$ (cfr. Eq. (11)), we also explored a situation with $\lambda=5.2$, which is usually encountered in experiments [51, 26]. In Fig. 5, $\text{D}_{\infty}$ (panel (a)) and $\theta_{\infty}$ (panel (b)) are reported as functions of Bq for the most confined simulated case (i.e., $\chi=0.7$) and with $\hat{h}=0.5$. In contract with $\text{D}_{\infty}$, which shows a perfect agreement between simulation data and Eq. (6), $\theta_{\infty}$ behaves as in Fig. 4, with a good deal with Eq. (11) for small values of Ca and a disagreement for $\text{Ca}=0.33$. This result further confirms the validity of our analytical predictions.
Conclusions – In this Letter, we fill the literature blanks concerning the estimation of steady-state behavior of a single drop under shear flow by presenting a unified analytical prediction describing the deformation (cfr. Eq. (6)) and inclination angle (cfr. Eq. (11)) of a confined/unconfined drop with/without interface viscosity. Since there are no available experiments for this system, to validate the quality and robustness of the proposed predictions, we performed immersed boundary-lattice Boltzmann simulations, which have already been validated in a series of previous works [38, 39, 41, 42, 40]. We also compared Eqs. (6) and (11) with the already known analytical predictions computed by Flumerfelt [35] for the unconfined drop with interface viscosity, showing that the accounting for confinement effects in Flumerfelt’s prediction is fundamental to quantitatively describe the correct deformation of a drop under these conditions. The robustness of our unified analytical predictions is strengthened by its application to different values of the drop viscosity ratio. We hope this result can open a new route to experiments of confined/unconfined drops with/without interface viscosity.
Acknowledgment – We thank L. Biferale and M. Sbragaglia for fruitful discussions. This work was supported by the Italian Ministry of University and Research (MUR) under the FARE program (No. R2045J8XAW), project “Smart-HEART”. FP acknowledges the support of the National Center for HPC, Big Data and Quantum Computing, Project CN_00000013 – CUP B93C22000620006, Mission 4 Component 2 Investment 1.4, funded by the European Union – NextGenerationEU.
@@
.1 Immersed boundary-lattice Boltzmann method and interface model
Instead of directly simulating the Navier-Stokes equations, the lattice Boltzmann (LB) method simulates the evolution of some probability distribution functions $\mbox{n}_{i}(\bm{\bm{x}},t)=f(\bm{\bm{x}},\bm{\bm{c}}_{i},t)$, representing the probability of finding fluid particles at position $\bm{\bm{x}}$ and time $t$, moving with a discrete velocity $\bm{\bm{c}}_{i}$. In particular, we implemented the so-called D3Q19 LB scheme, in which 19 discrete velocities are considered on a 3D regular lattice (see Ref. [54] for details). The dynamics of such functions $\mbox{n}_{i}$ is given by the LB equation:
$\mbox{n}_{i}^{\mbox{\tiny eq}}(\bm{\bm{x}},t)=w_{i}\rho\left(1+\frac{\bm{\bm{u}}\cdot\bm{\bm{c}}_{i}}{c_{s}^{2}}+\frac{(\bm{\bm{u}}\cdot\bm{\bm{c}}_{i})^{2}}{2c_{s}^{4}}-\frac{\bm{\bm{u}}\cdot\bm{\bm{u}}}{c_{s}^{2}}\right)\ ,$ (13)
Notice that $\tau_{\mbox{\tiny LB}}$ provides the link between the kinetic approach of the LB method and the hydrodynamic world of the Navier-Stokes equation via the fluid viscosity $\mu=\frac{\rho}{c_{s}^{2}}\left(\tau_{\mbox{\tiny LB}}-\frac{\Delta t}{2}\right)\frac{\Delta x^{2}}{\Delta t}$. Further technical details can be found in textbooks [54, 55] or in previous works by some of the authors, in which the same method (and the same numerical code) was employed [39, 48, 49, 41, 40, 42]. In the present work, the fluid domain is a three-dimensional rectangular box with sizes $\text{L}_{\mbox{\scriptsize x}}=\text{L}_{\mbox{\scriptsize y}}=128\ \Delta x$ and $\text{L}_{\mbox{\scriptsize z}}=H$, the latter varying to explore different degrees of confinement $\chi$. The immersed boundary method handles the two-way coupling between the fluid and the drop interface [54, 56]. The drop interface is modeled as a 3D triangular mesh with radius $R=20\ \Delta x$, made of $1.8\times 10^{4}$ faces. On each mesh Lagrangian node $k$ at position $\bm{\bm{r}}_{k}$, the interface force $\bm{\bm{\varphi}}_{k}(t)$ is computed (details on the interface model are given below), and then, according to the immersed boundary method, it is spread over the neighboring Eulerian (fluid, LB) nodes with coordinates $\bm{\bm{x}}$:
Concerning the interface model, we compute on each element of the triangular mesh the viscoelastic stress tensor $\bm{\bm{\tau}}=\bm{\bm{\tau}}_{s}+\bm{\bm{\tau}}_{v}$, that is made of the sum of the surface-tension stress tensor $\bm{\bm{\tau}}_{s}=\sigma\mathcal{I}_{2}$ [38, 39, 41, 42], with $\mathcal{I}_{2}$ being the 2D identity matrix, and the viscous tensor $\bm{\bm{\tau}}_{v}$, which is given by the Boussinesq-Scriven tensor [57]: $\bm{\bm{\tau}}_{v}=\mu_{\mbox{\scriptsize int}}[\nabla\bm{\bm{u}}+(\nabla\bm{\bm{u}})^{T}]$, where $\nabla\bm{\bm{u}}$ is the velocity gradient at the interface [38]. Numerical details on how to compute the nodal force $\bm{\bm{\varphi}}_{i}$ starting from the stress tensor $\bm{\bm{\tau}}$ can be found in Refs. [38, 39].
.2 Functions $f_{1}^{(\mbox{\tiny N})}$ and $f_{2}^{(\mbox{\tiny N})}$
@@
.3 Flumerfelt’s prediction for inclination angle

### v4

@@ IV Steady-state inclination angle
respectively. Since they considered an unconfined pure drop ($\chi=0$ and $\text{Bq}=0$), the functions $f_{i}$ have been chosen to recover the first-order expansion in Ca of the shape-evolution equation computed by Rallison [16]:
$f_{1}(\lambda)=f_{1}^{(\mbox{\tiny MM})}(\lambda)=\frac{40(\lambda+1)}{(2\lambda+3)(19\lambda+16)}\ ,$ (14a)
$f_{2}(\lambda)=f_{2}^{(\mbox{\tiny MM})}(\lambda)=\frac{5}{2\lambda+3}\ .$ (14b)
Since Rallison’s result has been recently extended by Narsimhan (N) [10] to include the effect of the interface viscosity, we matched the functions $f_{i}$ with the first-order expansion of the shape-evolution equation computed in Ref. [10], thus obtaining $f_{i}(\lambda,\text{Bq})=f_{i}^{(\mbox{\tiny N})}(\lambda,\text{Bq})$ (cfr. Eqs. (22) in Appendix .1). Notice that, in the limit of pure drop, $f_{i}^{(\mbox{\tiny N})}(\lambda,\text{Bq}=0)$ coincide with $f_{i}^{(\mbox{\tiny MM})}(\lambda)$. To include the effect of the confinement degree, we followed Minale’s approach [59], who extended $f_{i}^{(\mbox{\tiny MM})}$ to account for the effect of $\chi$. Thus, we extended $f_{i}$ as
$f_{1}^{\mathrm{ext}}=f_{1}^{(\mbox{\tiny N})}(\lambda,\text{Bq})/\psi_{1}(\lambda,\hat{h},\chi)\ ,$ (15a)
$f_{2}^{\mathrm{ext}}=f_{2}^{(\mbox{\tiny N})}(\lambda,\text{Bq})\psi_{2}(\lambda,\hat{h},\chi)\ ,$ (15b)
with
$\psi_{i}=1+C_{s}(\hat{h})\frac{\chi^{3}}{8}\tilde{f}_{i}\ ,$ (16)
where the expressions for $\tilde{f}_{i}$ have to be matched with the analytical results by Shapira & Haber [24] (in Ref. [59], $f_{i}^{(\mbox{\tiny N})}$ are replaced by $f_{i}^{(\mbox{\tiny MM})}$ in the definition of $f_{i}^{\mathrm{ext}}$). To obtain an expression for $\tilde{f}_{i}$, we substituted $f_{i}$ with $f_{i}^{\mathrm{ext}}$ in Eq. (13a) and, by neglecting $\mathcal{O}(\chi^{6})$, we obtained:
@@
We then considered the first-order expansion in Ca of Eq. (12) with the assumption $\Psi=\Psi_{\mathrm{SH}}$, i.e.,
$\mbox{D}(\lambda,\hat{h},\chi,\text{Ca},\text{Bq})=\text{Ca}\frac{19\lambda+16+32\text{Bq}}{16\lambda+16+32\text{Bq}}\left[1+C_{s}(\hat{h})\frac{\chi^{3}}{8}\frac{1+2.5\lambda}{1+\lambda}\right]\ .$ (18)
The assumption that $\Psi=\Psi_{\mathrm{SH}}$ follows from Shapira & Haber’s calculations [24]: indeed, the last hold for small values of $\chi$ and, as we showed in Fig. 2, $\Psi\approx\Psi_{\mathrm{SH}}$ for small degrees of confinement (i.e., $\chi<0.7$).
By comparing Eq. (17) with Eq. (18) it is straightforward to see that
$\tilde{f}_{1}+\tilde{f}_{2}=\frac{1+2.5\lambda}{1+\lambda}\ .$ (19)
This implies that, once $\tilde{f}_{2}$ is determined, $\tilde{f}_{1}$ is fixed. Since $\Psi_{{\color[rgb]{0,0,0}\mathrm{SH}}}(\lambda,\hat{h},\chi)$ (i.e., the term in squared parenthesis in Eq. (18)) does not depend on Bq, then neither $\tilde{f}_{1}$ nor $\tilde{f}_{2}$ depend on Bq. Therefore, the values computed by Minale [59] for a pure confined drop are still valid:
@@
All results hitherto discussed refer to the case of $\lambda=1$. To prove the robustness of the analytical expressions for both D (cfr. Eq. (12)) and $\theta$ (cfr. Eq. (21)), we also explored a situation with $\lambda=5.2$, which is usually encountered in experiments [59, 26]. In Fig. 5, $\text{D}_{\infty}$ (panel (a)) and $\theta_{\infty}$ (panel (b)) are reported as functions of Bq for the most confined simulated case (i.e., $\chi=0.7$) and with $\hat{h}=0.5$. In contrast with $\text{D}_{\infty}$, which shows a perfect agreement between simulation data and Eq. (12), $\theta_{\infty}$ behaves as in Fig. 4, with a good deal with Eq. (21) for small values of Ca and a disagreement for $\text{Ca}=0.33$. This result further confirms the validity of our analytical predictions.
V Conclusions
In this work, we present a comprehensive analytical prediction describing the steady-state deformation (cfr. Eq. (12)) and inclination angle (cfr. Eq. (21)) of a confined/unconfined drop with/without interface viscosity, in the limit of small deformation and negligible inertia effects. Since there are no available experiments for this system, to validate the quality and robustness of the proposed predictions, we performed immersed boundary-lattice Boltzmann simulations, which have already been validated in a series of previous works [38, 39, 41, 42, 40]. We also compared Eqs. (12) and (21) with the already known analytical predictions computed by Flumerfelt [35] for the unconfined drop with interface viscosity, showing that the accounting for confinement effects in Flumerfelt’s prediction is fundamental to quantitatively describe the correct deformation of a drop under these conditions. The robustness of our analytical predictions is strengthened by its application to different values of the drop viscosity ratio. We hope this result can open a new route to experiments of confined/unconfined drops with/without interface viscosity.
VI Acknowledgment
We thank L. Biferale and M. Sbragaglia for fruitful discussions. This work was supported by the Italian Ministry of University and Research (MUR) under the FARE program (No. R2045J8XAW), project “Smart-HEART”. FP acknowledges the support of the National Center for HPC, Big Data and Quantum Computing, Project CN_00000013 – CUP B93C22000620006, Mission 4 Component 2 Investment 1.4, funded by the European Union – NextGenerationEU.
@@ VI Acknowledgment
.1 Functions $f_{1}^{(\mbox{\tiny N})}$ and $f_{2}^{(\mbox{\tiny N})}$
@@
.2 Flumerfelt’s prediction for inclination angle